OpenStudy (anonymous):

Given the proportion a/b=6/20, what ratio completes the equivalent proportion a/6=-?

ganeshie8 (ganeshie8):

\(\large \frac{a}{b} = \frac{6}{20}\)

ganeshie8 (ganeshie8):

they just switched the 6 and b, you can think of it like when u cross multiply, 6 and b go together.

ganeshie8 (ganeshie8):

\(\large \frac{a}{b} = \frac{6}{20}\) then, \(\large \frac{a}{6} = \frac{b}{20}\)

OpenStudy (anonymous):

Is 6/20 the answer?

ganeshie8 (ganeshie8):

no. b/20 is the answer